OpenStudy (anonymous):

HELP ME PLEASEEEEEE!!!!!!!!!! (x+20) over 2 = 3x

Parth (parthkohli):

http://openstudy.com/study#/updates/50169891e4b04dfc808ab9f5 You posted this question again.

OpenStudy (anonymous):

because i didnt understand what you were doing

Parth (parthkohli):

You could have asked what I was doing, or have bumped the question to the top.

OpenStudy (anonymous):

im new to this thing

Parth (parthkohli):

Can you tell me what \[\left ( x + 20 \over 2\right ) \times 2 \] is?

OpenStudy (anonymous):

40 and 4

Parth (parthkohli):

\[ {x + 20 \over 2} \times 2 \Longrightarrow {x + 20 \over 2} \times {2 \over 1} \]Multiply the numerators and denominators.

OpenStudy (anonymous):

you lost me

Parth (parthkohli):

You have a rule \[ {a \over c} \times {b \over d} = {a \times b \over c \times d} \]

Parth (parthkohli):

Do you get it now? You multiply both numerators and denominators.

OpenStudy (anonymous):

no you lost me

Parth (parthkohli):

\[ {x + 20 \over 2} \times {2 } \Longrightarrow {x + 20 \over 2} \times {2\over 1} \Longrightarrow {(x + 20) \times 2 \over 2 \times 1} \] Where are you confused?